y. Solve the absolute value equation. [2x – 3) = |4x + 9/ SAMSUNG
Answer #1

g solve the absolute value equation 12x-3) = 14x +9) solution- Square both side (2x-3)2 = (4x +9)? 4x² 12x + 9 = 16x²+72x+81